# Multi-scale Self-similar Finite-time Blowups of the Constantin–Lax–Majda Model for the 3D Euler Equations

### v2

@@ Abstract.
We construct a new class of asymptotically self-similar finite-time blowups that have two collapsing spatial scales for the 1D Constantin–Lax–Majda model. The larger spatial scale measures the decreasing distance between the bulk of the solution and the eventual blowup point, while the smaller scale measures the shrinking size of the bulk of the solution. Similar multi-scale blowup phenomena have recently been discovered for many higher dimensional equations. Our study may provide some understanding of the common mechanism behind these multi-scale blowups.
${}^{1}$ dhuang@math.pku.edu.cn
School of Mathematical Sciences, Peking University
Dedicated to Thomas Yizhao Hou on the occasion of his 60th birthday.
@@ 1. Introduction
where $\Omega$ is referred to as the self-similar profile, $T>0$ is some finite blowup time, and $c_{\omega},c_{l}$ are the blowup powers for the magnitude and the spatial scaling, respectively. In particular, in the construction of Elgindi-Jeong, $\Omega(x)=-2x/(1+x^{2})$ (up to rescaling), $c_{\omega}=-1$, and $c_{l}=1$. Later, Elgindi–Ghoul–Masmoudi [12] improved on this result by establishing the stability of such self-similar blowup. Roughly speaking, solutions that are sufficiently close to an exact profile will also blowup in an asymptotic self-similar way as $t\rightarrow T-0$:
$\omega(x,t)=(T-t)^{c_{\omega}}\left(\Omega\left(\frac{x}{(T-t)^{c_{l}}}\right)+o(1)\right),$ (1.3)
@@
where the larger spatial scale $(T-t)^{c_{s}}$ (with a smaller power $c_{s}>0$) measures the distance between the origin and the center location of the bulk part of the solution, while the smaller spatial scale $(T-t)^{c_{l}}$ (with a larger power $c_{l}>0$) measures the size of the bulk part of the solution. The function $r(t)$ is continuously differentiable in $t$ and satisfies $0<\lim_{t\rightarrow T-0}r(t)=r(T)<+\infty$; it captures the precise location of the bulk part in the larger scale. That is, the peak location of the solution will converge to the origin like $r(t)(T-t)^{c_{s}}$, in the mean while the width of the peak will shrink even faster like $(T-t)^{c_{l}}$. It can also be understood as a traveling wave solution that collapses into one point and blows up in finite time. Note that a solution of (1.1) satisfies the scaling and translation property
@@
for any $\alpha,\gamma\in\mathbb{R}$ and $\beta>0$. Hence, one can always rescale the solution so that $T=1$ and $r(T)=1$ (and $\Omega$ will change accordingly).
In particular, we will use the explicit solution formula provided in [9] to construct two-scale self-similar blowups of the form (1.4) with
@@
for any positive integer $n\geq 1$ (Theorem 2.3). Note that in our construction $|c_{\omega}|$ and $c_{l}$ can be arbitrarily larger, while $c_{s}$ is always $1/2$. Again, suppose that $x=0$ is the only element of the set $\mathrm{S}$. We will show that such a two-scale blowup will happen around $x=0$ given additionally that $\omega_{0}^{(k)}(0)=0$, $k=1,2,\dots,2n$, and $\omega_{0}^{(2n+1)}(0)<0$. In other words, in contrast to the one-scale blowup case, the initial solution $\omega_{0}$ needs to be more degenerate at the blowup point for a two-scale blowup to happen.
@@
To better understand the multi-scale blowup phenomena in a uniform framework, we will perform a complex analysis that has been commonly used in the literature (e.g. [9, 13, 22]) for studying the CLM model (1.1). By considering the complex function $\eta(z,t)=\omega(z,t)+\mathrm{i}\bm{H}(\omega)(z,t)$ that solves a local ODE $\eta_{t}=-\mathrm{i}\eta^{2}/2$, we interpret the finite-time blowup of $\eta(z,t)$ on the real line $\mathbb{R}$ as one or several of its poles touch the real line for the first time. We will construct representative examples of $\eta$ whose poles all start in the lower half of the complex space $\mathbb{C}$ but then travel to the real line with different types of trajectories. Depending on the asymptotic trajectories of the moving poles, the restriction of $\eta$ on $\mathbb{R}$ will exhibit self-similar finite-time blowups with different scale features.
@@
Finally, we remark that multi-scale finite-time blowups have also been found in other types of PDE systems. For instance, Collot–Ghoul–Masmoudi–Nguyen [3] recently discovered collapsing-ring finite-time blowup solutions for the $d$-dimensional ($d\geq 3$) Keller–Segel system that have a very similar two-scale structure as in (1.4). In their construction, the mass concentrates near a thin sphere shell whose radius shrinks like $(T-t)^{1/d}$ while whose thickness shrinks like $(T-t)^{1-1/d}$.
Acknowledgement
@@ 2. Multi-scale blowups of the CLM models
$\begin{split}\omega(x,t)&=\frac{4\omega_{0}(x)}{(2-t\bm{H}(\omega_{0})(x))^{2}+t^{2}\omega_{0}(x)^{2}}\\ \bm{H}(\omega)(x,t)&=\frac{2\bm{H}(\omega_{0})(x)(2-t\bm{H}(\omega_{0})(x))-2t\omega_{0}(x)^{2}}{(2-t\bm{H}(\omega_{0})(x))^{2}+t^{2}\omega_{0}(x)^{2}},\end{split}$ (2.1)
exactly solves (1.1). As stated in [9, Corollary 1], it follows immediately from this formula that the solution $\omega(x,t)$ will blow up at a finite-time $T$ if and only if the set
@@
In what follows, we simply consider the case where $x=0$ is the only element of $\mathrm{S}$, i.e. $\mathrm{S}=\{0\}$, and so it must be the blowup point. We can ensure this by imposing odd symmetry and a sign condition on $\omega_{0}$: $\omega_{0}(-x)=-\omega_{0}(x)$, $\omega_{0}(x)<0$ for $x>0$. As a consequence, $\bm{H}(\omega_{0})(x)$ is an even function of $x$, and $\bm{H}(\omega_{0})(0)>0$. Note that these properties are preserved by the evolution of (1.1). We will argue that, the solution $\omega(x,t)$ will develop different types of self-similar blowups focusing at $x=0$, with different order of degeneracy of $\omega_{0}$ at $x=0$.
@@ 2.2. Two-scale self-similar blowup: Basic case
$\bm{H}(\omega_{0})(0)=-\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\omega_{0}(y)}{y}\,\mathrm{d}{y}>0,$
$\bm{H}(\omega_{0})^{\prime\prime}(0)=-\frac{2}{\pi}\int_{-\infty}^{\infty}\frac{\omega_{0}(y)}{y^{3}}\,\mathrm{d}{y}>0.$
@@
What Theorem 2.2 describes is a focusing finite-time blowup solution that has two spatial scales. The larger scale $(T-t)^{1/2}$ captures the traveling maximum location of the solution, while the smaller scale $T-t$ measures the characteristic width of the bulk of the solution. It can be viewed as a traveling wave that collapses into the origin and blows up in finite time. An example of this type of two-scale asymptotically self-similar blowup with initial data $\omega_{0}(x)=-2x^{3}/(4+x^{4})$ is presented in Figure 2.2.
@@ 2.3. Two-scale self-similar blowup: General case
$p(x):=\frac{\omega_{0}(x)}{-4cx^{2n+1}},\quad q(x):=\frac{H(\omega_{0})(x)-H(\omega_{0})(0)}{2bx^{2}}=\frac{1}{2b}\bm{H}\left(\frac{\omega_{0}}{x^{2}}\right)(x).$
@@
Since $q\in C^{2}(-\delta,\delta)$, $q(0)=1$, $q^{\prime}(0)=0$, we have $X(s)/s=q(X(s))^{-1/2}\in C^{1}([-s_{0},s_{0}])$, and
@@
The limits above imply that $g(\tau)\in C^{1}([0,s_{0}^{1/2}])$ where $g(\tau):=X(\tau^{1/2})/\tau^{1/2}$. In particular, we have
@@
Let us say a little more about traveling wave solutions. Suppose that $\omega_{0}(x)=f(x)$ is sufficiently smooth, decays sufficiently fast in the far-field, and does not have any zero on $\mathbb{R}$. If we require $\omega(x,t)=f(x+rt)$, $r\neq 0$, to be a solution of (1.1), then $f$ solves
@@
$\frac{\mathrm{d}{\,}}{\mathrm{d}{t}}\bm{H}(\omega)(0,t)=\frac{1}{2}\left(\bm{H}(\omega)(0,t)^{2}-\omega(0,t)^{2}\right)=\frac{1}{2}\bm{H}(\omega)(0,t)^{2}.$
implying that
$\bm{H}(\omega)(0,t)=\frac{2}{T-t},$
where $T=2/\bm{H}(\omega)(0,0)$. This is probably the simplest way to see the blowup nature of the CLM model. We have thus found a conserved quantity of the solution:
@@ 3. A complex analysis
$\begin{split}\omega_{t}&=\omega\bm{H}(\omega).\\ \bm{H}(\omega)_{t}&=\frac{1}{2}\bm{H}(\omega)^{2}-\frac{1}{2}\omega^{2}.\end{split}$
Consider the complex valued function $\eta(x,t):=\omega(x,t)+\mathrm{i}\bm{H}(\omega)(x,t)$. Then the equation of $\eta(x,t)$ is simply
@@
which is exactly the CLM model (1.1). The poles of $\eta(x,t)$, which initially lie in the interior of $\mathbb{C}_{-}$, may travel upward and touch the real line $\mathbb{R}$ at some finite time $T$. Then a singularity of $\omega$ will occur on $\mathbb{R}$ at time $T$. Hence, we can understand the blowup of $\omega$ by studying the dynamics of the poles of $\eta$.
@@
$\zeta_{t}=\frac{\mathrm{i}}{2},\quad\zeta(x,0)=\zeta_{0}(x)=1/\eta_{0}(x),$

### v3

@@ Abstract.
We construct a new class of asymptotically self-similar finite-time blowups that have two collapsing spatial scales for the 1D Constantin–Lax–Majda model. The larger spatial scale measures the decreasing distance between the bulk of the solution and the eventual blowup point, while the smaller scale measures the shrinking size of the bulk of the solution. Our construction is based on an asymptotic analysis applied to the explicit solution formula of the model, and a complex analysis is performed to provide an understanding of multi-scale blowups from a pole dynamics perspective. Similar multi-scale blowup phenomena have recently been discovered for many higher dimensional equations. Our study may provide some understanding of the common mechanism behind these multi-scale blowups.
1 School of Mathematical Sciences, Peking University. Email: dhuang@math.pku.edu.cn
2 Computing and Mathematical Sciences, California Institute of Technology. Email: xqin2@caltech.edu
3 School of Mathematical Sciences, Peking University. Email: wangxiuyuan@stu.pku.edu.cn
@@ 1. Introduction
where $\Omega$ is referred to as the self-similar profile, $T>0$ is some finite blowup time, and $c_{\omega},c_{l}$ are the blowup powers for the magnitude and the spatial scaling, respectively. In particular, in the construction of Elgindi–Jeong, $\Omega(x)=-2x/(1+x^{2})$ (up to rescaling), $c_{\omega}=-1$, and $c_{l}=1$. Later, Elgindi–Ghoul–Masmoudi [16] improved on this result by establishing the stability of such self-similar blowup. Roughly speaking, solutions that are sufficiently close to an exact profile will also blowup in an asymptotically self-similar way as $t\rightarrow T-0$:
$\omega(x,t)=(T-t)^{c_{\omega}}\left(\Omega\left(\frac{x}{(T-t)^{c_{l}}}\right)+o(1)\right).$ (1.3)
@@
where the larger spatial scale $(T-t)^{c_{s}}$ (with a smaller power $c_{s}>0$) measures the distance between the origin and the center location of the bulk part of the solution, while the smaller spatial scale $(T-t)^{c_{l}}$ (with a larger power $c_{l}>0$) measures the size of the bulk part of the solution. The function $r(t)$ is continuously differentiable in $t$ and satisfies $0<\lim_{t\rightarrow T-0}r(t)=r(T)<+\infty$; it captures the precise location of the bulk part in the larger scale. That is, the peak location of the solution will converge to the origin like $r(t)(T-t)^{c_{s}}$, in the meanwhile the width of the peak will shrink even faster like $(T-t)^{c_{l}}$. It can also be understood as a traveling wave solution that collapses into one point and blows up in finite time. In fact, we will show that the self-similar profile $\Omega$ in the two-scale blowup (1.4) is an exact traveling wave solution to the CLM model (1.1). Note that a solution of (1.1) satisfies the scaling and translation property
@@
for any $\alpha,\gamma\in\mathbb{R}$ and $\beta>0$. Hence, one can always rescale the solution so that $T=1$ and $r(T)=1$ (and $\Omega$ will change accordingly). In particular, we will use the explicit solution formula provided in [13] to construct two-scale self-similar blowups of the form (1.4) with
@@
for any positive integer $n\geq 1$ (Theorem 2.3). Note that in our construction $|c_{\omega}|$ and $c_{l}$ can be arbitrarily large, while $c_{s}$ is always $1/2$. Again, suppose that $x=0$ is the only element of the set $\mathrm{S}$. We will show that such a two-scale blowup will happen around $x=0$ given additionally that $\omega_{0}^{(k)}(0)=0$, $k=1,2,\dots,2n$, and $\omega_{0}^{(2n+1)}(0)<0$. In other words, in contrast to the one-scale blowup case, the initial solution $\omega_{0}$ needs to be more degenerate at the blowup point for a two-scale blowup to happen.
@@
To better understand the multi-scale blowup phenomena in a uniform framework, we will perform a complex pole dynamics analysis that has been commonly used in the literature (e.g. [13, 17, 29, 1, 32]) for studying the CLM model (1.1). By considering the complex function $\eta(z,t)=\omega(z,t)+\mathrm{i}\bm{H}(\omega)(z,t)$ that solves a local ODE $\eta_{t}=-\mathrm{i}\eta^{2}/2$, we interpret the finite-time blowup of $\eta(z,t)$ on the real line $\mathbb{R}$ as one or several of its poles touch the real line for the first time. We will construct representative examples of $\eta$ whose poles all start in the lower half of the complex space $\mathbb{C}$ but then travel to the real line with different types of trajectories. Depending on the asymptotic trajectories of the moving poles, the restriction of $\eta$ on $\mathbb{R}$ will exhibit self-similar finite-time blowups with different scale features.
@@
We also want to mention a two-scale phenomenon that appears in the head-on collision of two anti-parallel vortex rings in 3D axisymmetric flows without swirl. In the studies of Childress [10] and Childress–Gilbert–Valiant [5], numerical observations and formal calculations suggest that the vorticity concentrated in the vortex rings grows as $t^{4/3}$ (an infinite-time blowup), where the blowup rate $4/3$ was recently proved to be optimal in such scenario [28]. In the meantime, the radius of each vortex ring grows like $t^{4/3}$, while the cross-sectional area of each ring shrinks like $t^{-2}$, demonstrating a spatial two-scale feature in this blowup process. Remarkably, it is observed that, as time goes to infinity, the cross-sectional region converges to the shape of the Sadovskii vortex-patch dipole [31], which is a steady traveling wave solution to the 2D incompressible Euler equation. It is worth mentioning that the theoretical existence of the Sadovskii vortex-patch dipole was only recently proved by two groups, Huang–Tong [24] and Choi–Jeong–Sim [12], using two completely different approaches (one based on a fixed-point method and the other based on a variational method).
@@
Finally, we remark that multi-scale finite-time blowups have also been found in other types of PDE systems. For instance, Collot–Ghoul–Masmoudi–Nguyen [4] recently discovered collapsing-ring finite-time blowup solutions for the $d$-dimensional ($d\geq 3$) Keller–Segel system that have a very similar two-scale structure as in (1.4). In their construction, the mass concentrates near a thin sphere shell whose radius shrinks like $(T-t)^{1/d}$ while whose thickness shrinks like $(T-t)^{1-1/d}$ ($T$ is the finite blowup time).
@@ 2. Multi-scale blowups of the CLM models
$\begin{split}\omega(x,t)&=\frac{4\omega_{0}(x)}{(2-t\bm{H}(\omega_{0})(x))^{2}+t^{2}\omega_{0}(x)^{2}},\\ \bm{H}(\omega)(x,t)&=\frac{2\bm{H}(\omega_{0})(x)(2-t\bm{H}(\omega_{0})(x))-2t\omega_{0}(x)^{2}}{(2-t\bm{H}(\omega_{0})(x))^{2}+t^{2}\omega_{0}(x)^{2}},\end{split}$ (2.1)
exactly solves (1.1). As stated in [13, Corollary 1], it follows immediately from this formula that the solution $\omega(x,t)$ will blow up at a finite-time $T$ if and only if the set
@@
In what follows, we simply consider the case where $x=0$ is the only element of $\mathrm{S}$, i.e. $\mathrm{S}=\{0\}$, and so it must be the blowup point. We can ensure this by imposing odd symmetry and a sign condition on $\omega_{0}$: $\omega_{0}(-x)=-\omega_{0}(x)$, $\omega_{0}(x)<0$ for $x>0$. As a consequence, $\bm{H}(\omega_{0})(x)$ is an even function of $x$, and $\bm{H}(\omega_{0})(0)>0$. Note that these properties are preserved by the evolution of (1.1). We will argue that, the solution $\omega(x,t)$ will develop different types of self-similar blowups focusing at $x=0$, with different orders of degeneracy of $\omega_{0}$ at $x=0$.
@@ 2.2. Two-scale self-similar blowup: Basic case
What Theorem 2.2 describes is a focusing finite-time blowup solution that has two spatial scales. The larger scale $(T-t)^{1/2}$ captures the traveling maximum location of the solution, while the smaller scale $T-t$ measures the characteristic width of the bulk part of the solution. It can be viewed as a traveling wave solution that collapses into the origin and blows up in finite time. An example of this type of two-scale asymptotically self-similar blowup with initial data $\omega_{0}(x)=-2x^{3}/(4+x^{4})$ is demonstrated in Figure 2.2.
@@ 2.3. Two-scale self-similar blowup: General case
$p(x):=\frac{\omega_{0}(x)}{-4cx^{2n+1}},\quad q(x):=\frac{\bm{H}(\omega_{0})(x)-\bm{H}(\omega_{0})(0)}{2bx^{2}}=\frac{1}{2b}\bm{H}\left(\frac{\omega_{0}}{x^{2}}\right)(x).$ (2.4)
@@
Note that, since $q\in C^{2}(-\delta,\delta)$, $q(0)=1$, $q^{\prime}(0)=0$, we have $X(s)/s=q(X(s))^{-1/2}\in C^{1}([-s_{0},s_{0}])$, and
@@
The limits above imply that $g(\tau)\in C^{1}([0,s_{0}^{2}])$ where $g(\tau):=X(\tau^{1/2})/\tau^{1/2}$. In particular, we have
@@ 2.4. Relation to traveling wave solutions
Let us say a little more about traveling wave solutions. Suppose that $\omega_{0}(x)=f(x)$ is sufficiently smooth, decays sufficiently fast in the far field, and does not have any zero on $\mathbb{R}$. If we require $\omega(x,t)=f(x+rt)$, $r\neq 0$, to be a solution of (1.1), then $f$ solves
@@
$\frac{\mathrm{d}{\,}}{\mathrm{d}{t}}\bm{H}(\omega)(0,t)=\frac{1}{2}\left(\bm{H}(\omega)(0,t)^{2}-\omega(0,t)^{2}\right)=\frac{1}{2}\bm{H}(\omega)(0,t)^{2},$
which implies $\bm{H}(\omega)(0,t)=2(T-t)^{-1}$ with $T=2/\bm{H}(\omega)(0,0)$. This is probably the simplest way to see the blowup nature of the CLM model. We have thus found a conserved quantity of the solution:
@@
2.5. Non-smooth initial data
Other than the one-scale blowup (1.2) with a smooth profile function $\Omega$, Elgindi–Jeong [17] also constructed one-scale self-similar blowup solutions with Hölder continuous profiles for the CLM model. More specifically, they found a family of self-similar solutions of the form
$\omega(x,t)=(T-t)^{-1}\cdot\Omega_{\alpha}\left(\frac{x}{(T-t)^{1/\alpha}}\right),$ (2.15)
with $\alpha\in(0,1)$, where the profile (up to rescaling) is a one-point $C^{\alpha}$ function given by
$\Omega_{\alpha}(x)=\frac{-\sin(\alpha\pi/2)\operatorname{sgn}{(}x)|x|^{\alpha}}{1+2\cos(\alpha\pi/2)|x|^{\alpha}+|x|^{2\alpha}},$ (2.16)
which is odd in $x$ and smooth at any $x\neq 0$. Stability of this family of Hölder continuous profiles was also later established in [16, 8], and hence an asymptotically self-similar blowup of the form (1.3) can happen for odd initial data $\omega_{0}$ such that $\omega_{0}(x)\sim-\operatorname{sgn}{(}x)|x|^{\alpha}$ for $x\sim 0$.
Therefore, it is also interesting to consider two-scale self-similar blowup of the CLM model with initial data that are degenerate but less regular at the origin. In particular, we may consider
$\omega_{0}(x)=-\operatorname{sgn}{(}x)|x|^{\beta}p(x)$ (2.17)
for some $\beta>1$ that is not an odd integer and for some smooth positive even function $p$ that decays sufficiently fast in the far field. One can again define $q(x)$ as in (2.4). When $\beta>4$, it is straightforward to check by the definition of the Hilbert transform that $q$ is at least $C^{2}$ near the origin and the expansion $q(x)=q(0)+O(x^{2})$ still holds. Then, we can simply re-implement the proof of Theorem 2.3 (but omit the tedious details here) to show that as $t\rightarrow T-0$, for $x=r(t)(T-t)^{1/2}+O((T-t)^{(\beta-1)/2})$,
$\omega(x,t)=\frac{1}{(T-t)^{\beta/2}}\left(\Omega_{2}\left(\frac{x-r(t)(T-t)^{1/2}}{(T-t)^{(\beta-1)/2}}\right)+o(1)\right),$ (2.18)
where $T$ and $r(t)$ are given as in Theorem 2.3, and $\Omega_{2}$ is again a traveling wave solution to (1.1) that can be rescaled to $-1/(1+x^{2})$. When $\beta\in(2,3)\cup(3,4]$, the corresponding function $q$ is less regular: for $x\sim 0$, $q(x)=q(0)+O(|x|^{\beta-2})$ with $\beta\in(2,3)\cup(3,4)$ and $q(x)=q(0)+O(|x|^{2}\ln|x|)$ with $\beta=4$. Nevertheless, one can slightly modify our proof and still show that the asymptotic expression (2.18) is valid for $x=r(t)(T-t)^{1/2}+O((T-t)^{(\beta-1)/2})$. Therefore, a two-scale self-similar finite-time blowup also happens with initial data of the form (2.17) for $\beta>2$.
The case $\beta=2$ is a bit more complicated but also interesting by itself, as in this case $q(x)\sim\ln(1/|x|)$ for $x\sim 0$. With a similar strategy, one can substitute the assumption (2.17) into the explicit solution formula (2.1) to prove that, for $x=r(t)\big{(}(T-t)^{1/2}/|\ln(T-t)|^{1/2}\big{)}+O\big{(}(T-t)^{1/2}/|\ln(T-t)|^{3/2}\big{)}$,
$\omega(x,t)=\frac{|\ln(T-t)|}{T-t}\left(\Omega_{2}\left(\frac{x-r(t)\big{(}(T-t)^{1/2}/|\ln(T-t)|^{1/2}\big{)}}{(T-t)^{1/2}/|\ln(T-t)|^{3/2}}\right)+o(1)\right),$
for some $T$ and $r(t)$ that can be determined in the proof. That is, a two-scale self-similar finite-time blowup still happens, but the ratio between the larger scale and the smaller scale is only $|\ln(T-t)|$.
As for $\beta\in(1,2)$, one should expect a one-scale self-similar blowup to happen. In fact, the non-smooth function (2.16) constructed in [17] is also an exact one-scale self-similar profile for $\alpha=\beta\in(1,2)$.
2.6. On the blowup rates
@@ 3.1. Complexification of the model
$\begin{split}\omega_{t}&=\omega\bm{H}(\omega),\\ \bm{H}(\omega)_{t}&=\frac{1}{2}\bm{H}(\omega)^{2}-\frac{1}{2}\omega^{2}.\end{split}$
Consider the complex-valued function $\eta(x,t):=\omega(x,t)+\mathrm{i}\bm{H}(\omega)(x,t)$. Then the equation of $\eta(x,t)$ is simply
@@ 3.2. Pole dynamics
which is exactly the CLM model (1.1). The poles of $\eta(z,t)$, which initially lie in the interior of $\mathbb{C}_{-}$, may travel upward and touch the real line $\mathbb{R}$ at some finite time $T$. Then a singularity of $\omega$ will occur on $\mathbb{R}$ at time $T$. Hence, we can understand the blowup of $\omega$ by studying the dynamics of the poles of $\eta$.
@@
$\zeta_{t}=\frac{\mathrm{i}}{2},\quad\zeta(z,0)=\zeta_{0}(z)=1/\eta_{0}(z),$